# Your Causal Self-Attentive Recommender Hosts a Lonely Neighborhood

### v2

@@ Abstract.
ding (AE) and uni-directional/auto-regressive (AR) attention mechanisms, where the conclusions regarding architectural and performance superiority remain inconclusive. Previous efforts in such comparisons primarily involve summarizing existing works to identify a consensus or conducting ablation studies on peripheral modeling techniques, such as choices of loss functions. However, far fewer efforts have been made in (1) theoretical and (2) extensive empirical analysis of the self-attention module, the very pivotal structure on which performance and designing insights should be anchored. In this work, we first provide a comprehensive theoretical analysis of AE/AR attention matrix in the aspect of (1) sparse local inductive bias, a.k.a neighborhood effects, and (2) low rank approximation. Analytical metrics reveal that the AR attention exhibits sparse neighborhood effects suitable for generally sparse recommendation scenarios. Secondly, to support our theoretical analysis, we conduct extensive empirical experiments on comparing vanilla and variant AE/AR attention on five popular benchmarks with AR performing better overall. Results based on adaptive tuning, modularized design and Huggingface are reported. Lastly, we shed light on future design choices for performant self-attentive recommenders. We make our code and data available at https://github.com/yueqirex/Self-Attention-Direction-Check.
@@ 1. Introduction
AR-based and AE-based self-attentive models are widely used in sequential recommendation tasks (Liu et al., 2021b; Fan et al., 2021; Petrov and Makarov, 2021). AR-based recommenders, represented by SASRec (Kang and McAuley, 2018), SSE-PT (Wu et al., 2020), are commonly compared with AE-based recommenders, represented by BERT4Rec (Sun et al., 2019), LOCKER (He et al., 2021) in a wide range of recommendation datasets. However, those comparisons between AR-based and AE-based recommenders always lead to discrepancies in conclusions from a number of papers (Petrov and Macdonald, 2022). Specifically, while AE models claim superiority of bi-directional attention over AR models in (Ma et al., 2020; Zhou et al., 2021; Liu et al., 2021a; Li et al., 2021b; Cho et al., 2020; Zhang et al., 2020; Tong et al., 2021; Yue et al., 2021; Wu et al., 2021; Sun et al., 2022), AR models show better performance in (Zhou et al., 2020; Li et al., 2021a; Zhang et al., 2021; Qiu et al., 2022). Approaching this issue, (Petrov and Macdonald, 2022) focuses on replicating the original results of the AE-based recommender BERT4REC, rectifying several implementation details to obtain closer results to the reported. Afterwards, (Klenitskiy and Vasilev, 2023) shows that the AR-based recommender, SASRec, outperforms the AE-based BERT4REC when using the same cross-entropy loss function of the latter.
@@ 3.1. Smooth sparsity measurement of attention matrix
Following the convention proposed by (Su, 2023) and in matrix analysis, sparsity traditionally refers to the proportion of zero elements relative to the total. However, for attention matrices, the concept differs: sparsity signifies that most entries are negligible or near zero, with only a minority exhibiting substantial values. As such, conventional sparsity metrics fall short in capturing this nuance. Quantifying sparsity in attention matrices necessitates a clear definition of what constitutes a ”significant” value. A practical method involves computing the average absolute value of all elements, normalized by the maximum absolute value:
@@
The introduction of the max operation in Equation 2 compromises smoothness, rendering the measure highly susceptible to outliers. Leveraging approximation theory, the LogSumExp (LSE) function, defined in Equation 3, offers a smooth alternative to the maximum function. This inherent smoothness effectively attenuates the influence of extreme values within the attention matrix $X$.
@@
Observe that $(|x_{1}|+|x_{2}|+\ldots+|x_{n}|)/n$ corresponds to the $L_{1}$-norm of $\mathbf{x}$. Substituting the $L_{1}$-norm and Equation 4 into Equation 2 yields Equation 5. Empirically, $p$ usually takes the values of 2 and omitting the scaling factor of $n$ given fixed size vector/matrix comparison. Under these assumptions, the resulting sparsity measure exhibits several key properties, which we will explore in greater detail.
@@ 4.2. Approximation results, visualization and interpretability
Figure 2. Attention visualizations. First row is original matrix; Second row is corresponding low-rank approximations using top-5 largest singular values. Pure BERT4Rec-AE attention has a clear-pattern rank-5 approximation while local-attn and AR losses their patterns and need a higher-rank approximation.
@@
Figure 3. Singular values distribution for random user X in descending order for BERT4Rec-{AE, AR}, AE’s singular values drops to near zero in a faster rate than AR, suggesting a lower rank approximation.
@@ 5.2.1. Experimental setup
where SSE-PT (Wu et al., 2020) is an effective way to combine user embeddings and item embeddings without causing overfitting; Local-attn (He et al., 2021) apply a window-based local attention mechanism in addition to vanilla self-attention. {Loss-CE, Loss-BPR} serves as two mostly used loss functions for sequential recommendation; More details of comparison models are introduced in Section 2. Due to computational resource limitations, we can not study every combination of designs but rather set one representative configuration for each aspect in {Feature, Modeling, Optimization}. For each of above three designs, we search the best hyperparameters for AE and AR respectively and compare their performance. To tune as many combinations of hyper-parameters with limited machines, we adopt Ray tune and Asynchronous Successive Halving (ASHA) heuristics for hyper-parameter tuning shown in Figure 4. ASHA dynamically compares each configuration and early-terminates bad-performing ones to save resources. This way we derive more optimal hyper-parameters for each model given limited machines. Evaluation results for different design options are discussed in section 5.2.2. AE / AR models of each design choice are derived by adding/removing the causal mask in the attention module.
@@ 5.3.1. Experimental setup
Bridging the gap between the Sequential Recommenders and fast-developing NLP models has been long talked about (de Souza Pereira Moreira et al., 2021). Here, we focus on further extensive study of AE and AR comparison on the latest NLP models from HuggingFace. Specifically, we select three popular models that are originally uni-directional (Llama, Transformer-XL) or bi-directional (ALBERT). Similar to the approach discussed in Section 5.2, we change the original models to their AE/AR counterparts by (1) adding/removing the causal mask and (2) changing their optimization task to the next item prediction/Mask Language Model (MLM). We continue to utilize the adaptive search for hyperparameters introduced in Section 5.2.1. Evaluation results for HuggingFace models are discussed in Section 5.3.2.
@@ 5.3.2. Experimental results
In summary, AR continues to show superior performance than AE on advanced HuggingFace NLP models, this is consistent with our conclusions from Section 5.1 and Section 5.2. AR’s performance gain is more significant on sparse datasets and it shows better ranking ability than AE by having more improved NDCG than Recall.
@@ 6. Connecting empirical results with theoretical analysis
Figure 5. Average performance visualizations. Figure 5(a) shows the average performance indeed increases with growing local inductive bias consistent with our theoretical analysis. Local-attn in Figure 5(b) shows lowest improvements due to already injected inductive bias in Local-attn-AE.

### v3

@@ Abstract.
ding (AE) and uni-directional/auto-regressive (AR) attention mechanisms, where the conclusions regarding architectural and performance superiority remain inconclusive. Previous efforts in such comparisons primarily involve summarizing existing works to identify a consensus or conducting ablation studies on peripheral modeling techniques, such as choices of loss functions. However, far fewer efforts have been made in (1) theoretical and (2) extensive empirical analysis of the self-attention module, the very pivotal structure on which performance and designing insights should be anchored. In this work, we first provide a comprehensive theoretical analysis of AE/AR attention matrix in the aspect of (1) sparse local inductive bias, a.k.a neighborhood effects, and (2) low rank approximation. Analytical metrics reveal that the AR attention exhibits sparse neighborhood effects suitable for generally sparse recommendation scenarios. Secondly, to support our theoretical analysis, we conduct extensive empirical experiments on comparing AE/AR attention on five popular benchmarks with AR performing better overall. Empirical results reported are based on our experimental pipeline named Modularized Design Space for Self-Attentive Recommender (ModSAR), supporting adaptive hyperparameter tuning, modularized design space and Huggingface plug-ins. We invite the recommendation community to utilize/contribute to ModSAR to (1) conduct more module/model-level examining beyond AE/AR comparison and (2) accelerate state-of-the-art model design. Lastly, we shed light on future design choices for performant self-attentive recommenders. We make our pipeline implementation and data available at https://github.com/yueqirex/SAR-Check.
@@ 1. Introduction
AR-based and AE-based self-attentive models are widely used in sequential recommendation tasks (Liu et al., 2021b; Petrov and Makarov, 2021; Fan et al., 2021). AR-based recommenders, represented by SASRec (Kang and McAuley, 2018), SSE-PT (Wu et al., 2020), are commonly compared with AE-based recommenders, represented by BERT4Rec (Sun et al., 2019), LOCKER (He et al., 2021) in a wide range of recommendation datasets. However, those comparisons between AR-based and AE-based recommenders always lead to discrepancies in conclusions from a number of papers (Petrov and Macdonald, 2022). Specifically, while AE models claim superiority of bi-directional attention over AR models in (Ma et al., 2020; Zhou et al., 2021; Liu et al., 2021a; Li et al., 2021b; Cho et al., 2020; Zhang et al., 2020; Tong et al., 2021; Yue et al., 2021; Wu et al., 2021; Sun et al., 2022), AR models show better performance in (Zhou et al., 2020; Li et al., 2021a; Zhang et al., 2021; Qiu et al., 2022). Approaching this issue, (Petrov and Macdonald, 2022) focuses on replicating the original results of the AE-based recommender BERT4Rec, rectifying several implementation details to obtain closer results to the reported. Afterwards, (Klenitskiy and Vasilev, 2023) shows that the AR-based recommender, SASRec, outperforms the AE-based BERT4Rec when using the same cross-entropy loss function of the latter.
@@ 3.1. Smooth Sparsity Measurement of Attention Matrix
For the attention matrix, we first review (1) why traditional sparsity measurement falls short and (2) the mathematical derivation of a more scientifically adopted metric proposed by (Su, 2023). Sparsity traditionally (Hurley and Rickard, 2009) refers to the proportion of zero elements relative to the total. However, for attention matrices, the concept differs: sparsity signifies that most entries are negligible or near zero, with only a minority exhibiting substantial values. Quantifying sparsity in attention matrices necessitates a clear definition of what constitutes a ”significant” value. A practical method involves computing the average absolute value of all elements, normalized by the maximum absolute value:
@@
The introduction of the max operation in Equation 2 compromises smoothness, rendering the measure highly susceptible to outliers. Leveraging approximation theory, the LogSumExp (LSE) function, defined in Equation 3, offers a smooth alternative to the maximum function. Within the attention matrix $X$, this inherent smoothness effectively attenuates the influence of extreme values (Su, 2023).
@@
Observe that $(|x_{1}|+|x_{2}|+\ldots+|x_{n}|)/n$ corresponds to the $L_{1}$-norm of $\mathbf{x}$. Substituting the $L_{1}$-norm and Equation 4 into Equation 2 yields Equation 5. Empirically, $p$ usually takes the values of 2 and omitting the scaling factor of $n$ given fixed size vector/matrix comparison. Under these assumptions, the resulting sparsity measure (Su, 2023) exhibits several key properties, which we will explore in greater detail.
@@ 4.2. Approximation Results, Visualization and Interpretability
Figure 2. Attention visualizations. First row is original matrix; Second row is corresponding low-rank approximations using top-5 largest singular values. Vanilla-AE (BERT4Rec-like here) attention has a clear-pattern rank-5 approximation while local-attn and AR losses their patterns and need a higher-rank approximation.
@@
Figure 3. Singular values distribution for random user X in descending order for Vanilla-{AE, AR} (BERT4Rec-like here), AE’s singular values drops to near zero in a faster rate than AR, suggesting a lower rank approximation.
@@ 5.2.1. Experimental Setup
where SSE-PT (Wu et al., 2020) is an effective way to combine user embeddings and item embeddings without causing overfitting; Local-attn (He et al., 2021) apply a window-based local attention mechanism in addition to vanilla self-attention. {Loss-CE, Loss-BPR} serves as two mostly used loss functions for sequential recommendation; More details of comparison models are introduced in Section 2. Due to computational resource limitations, we can not study every combination of designs but rather set one representative configuration for each aspect in {Feature, Modeling, Optimization}. For each of above three designs, we search the best hyperparameters for AE and AR respectively and compare their performance. To tune as many combinations of hyper-parameters with limited machines, we adopt Ray tune and Asynchronous Successive Halving (ASHA) heuristics for hyper-parameter tuning. ASHA dynamically compares each configuration and early-terminates bad-performing ones to save resources. The entire experimental pipeline is named Modularized Design Space for Self-Attentive Recommender (ModSAR) shown in Figure 4. Evaluation results for different design options are discussed in section 5.2.2. AE / AR models of each design choice are derived by adding/removing the causal mask in the attention module.
@@ 5.3.1. Experimental Setup
Bridging the gap between the Sequential Recommenders and fast-developing NLP models has been long talked about (de Souza Pereira Moreira et al., 2021). Here, we focus on further extensive study of AE and AR comparison on the latest NLP models from HuggingFace. Specifically, we select three popular models that are originally uni-directional (Llama, Transformer-XL) or bi-directional (ALBERT). Similar to the approach discussed in Section 5.2, we change the original models to their AE/AR counterparts by (1) adding/removing the causal mask and (2) changing their optimization task to the next item prediction/Mask Language Model (MLM). We keep here the adaptive hyperparameter search introduced in Section 5.2.1. Evaluation results for HuggingFace models are discussed in Section 5.3.2.
@@ 5.3.2. Experimental Results
In summary, AR continues to show superior performance than AE on advanced HuggingFace NLP models, this is consistent with our conclusions from Section 5.1 and Section 5.2. AR’s performance gain is more significant on sparse datasets. AR shows better ranking ability than AE by having a more improved NDCG than Recall.
@@ 6. Connecting Empirical Results with Theoretical Analysis
Figure 5. Figure 5(a) shows the average performance increases with growing local inductive bias consistent with our theoretical analysis. Local-attn in Figure 5(b) shows lowest improvements due to already injected inductive bias in Local-attn-AE.